# On a generalisation of Cameron’s base size conjecture

### v2

@@ 4. Fixed point ratios for classical groups
where $\phi$ is a field automorphism of order 2, and so we compute
@@
Finally, if $x$ is a field automorphism of order $r$, then the bounds presented at the start of [16, Section 3.2] and [16, 2.3] are sufficient. ∎
@@
If $x$ is a field, graph, or graph-field automorphism, then the bounds in [15, 8.1] suffice, so it is left to check the claim for semisimple and unipotent elements. In those cases we follow the same arguments as in [15, 8.1]. We start by noting that $H\cap{\rm PGL}(V)\leqslant{\rm PGSp}_{8}(q)$, and so $|x^{G_{I}}\cap H|\leqslant|x^{{\rm PGSp}_{8}(q)}|$. Now the conjugacy classes in ${\rm PGSp}_{8}(q)$ and ${\rm PGL}_{8}^{\epsilon}(q)$ are discussed in Section 3.2 briefly and in [22, Chapter 7] in more detail, and the centraliser structure for each class can be read off from [22, Tables B.3 and B.7], so one can check that the claim holds. ∎
@@
    $\epsilon=+$, $x$ is semisimple with $C_{G}(x)$ of type ${\rm GL}_{3}(q)^{2}$, and $\mbox{{\rm fpr}}(x,G/H)<|x^{G}|^{-0.32}$
@@
    $\epsilon=$-, $x$ is semisimple with $C_{G}(x)$ of type ${\rm GL}_{3}(q^{2})$, and $\mbox{{\rm fpr}}(x,G/H)<|x^{G}|^{-0.32}$.
@@
Here we may assume that $H$ is of type ${\rm Sp}_{6}(q)$ since the claim follows directly by Theorem 3.9 in all other cases. First assume $x$ is semisimple of odd order and let $i$ be minimal such that $r|q^{i}-1$. Again, we may assume that $x\in H$ and in particular that $r$ divides $|H|$, since otherwise we have ${\rm fpr}(x,G/H)=0$. Thus, we have $i\in\{1,2,3,4,6\}$. The embedding here is straightforward so we will only give details for the cases where $i\in\{1,4\}$ and leave the remaining cases for the reader.
@@
Now suppose that $x$ is unipotent. By appealing to Section 3.2, we find that the Jordan decompositions corresponding to an element $x\in H$ are
@@
Finally, if $x$ is a graph field or graph field automorphism or a semisimple involution, then the bounds in the proof of [15, 8.1] suffice. ∎
@@
The claim can be easily verified using Magma for $q\in\{2,4\}$, and so we may assume that $q\geqslant 8$ for the remainder of the proof. By inspecting [12, Tables 8.28 and 8.29], we find that if $H$ is not of type $G_{2}(q)$, then $H_{0}=H\cap G_{0}\in\{{\rm Sp}_{2}(q)\wr S_{3},{\rm Sp}_{2}(q^{3}){:}3,{\rm Sp}_{6}(q_{0})\}$, where $k$ is a prime and $q=q_{0}^{k}$.
@@
Now suppose that $x\in{\rm PGL}(V)$. There are two cases to consider depending whether or not $x^{G}\cap(H\setminus B)\neq\emptyset$. If $x^{G}\cap(H\setminus B)\neq\emptyset$, then either $r=2$ or $r=3$. If $r=2$, then by inspecting the proof of [16, 2.8], we deduce that $\log|x^{G_{I}}\cap H|/\log|x^{G_{0}}|<0.543$, so we may now assume that $r=3$. In this case we may assume $x=b\pi=(x_{1},x_{2},x_{3})\pi$, where $b\in B$ and $\pi\in S_{3}$ is a $3$-cycle and it follows from the proof of [63, 4.5] that $x$ is $B$-conjugate to $\pi$. Now note that if $y=(A,B,C)\in C_{B}(x)$, then since $\pi$ permutes $A,B$, and $C$, we must have $A=B=C$, and so
@@
We now note that with respect to a symplectic basis $\{e_{1},e_{2},e_{3},f_{1},f_{2},f_{3}\}$, $(I_{2},I_{2},I_{2})\pi$ embeds in ${\rm Sp}_{6}(q)$ as the block-diagonal matrix $[A,A]$, where
$A=\begin{pmatrix}0&0&1\\ 1&0&0\\ 0&1&0\end{pmatrix}$
and since $r=3$ and $q$ is even, $x$ is semisimple. Moreover, since $q$ is even, $3$ must either divide q1q-1 or $q+1$, and so if $i$ is minimal such that $3$ divides $q^{i}-1$, we must have $i=1$ or $2$. Now note that $x$ has eigenvalues $1^{2},(\omega)^{2},(\omega^{2})^{2}$, where $\omega\in\mathbb{F}_{q^{i}}$ is a primitive cube root of unity, and so $C_{G}(x)$ is of type ${\rm GL}_{2}^{\epsilon}(q)\times{\rm Sp}_{2}(q)$, where $\epsilon=(-1)^{i+1}$. Therefore,
@@
We can now compute bounds for $|x^{G_{I}}\cap H|$ and $|x^{G_{0}}|$ using [22, Table B.7]. For example, if $C_{G}(x)$ is of type ${\rm GU}_{2}(q)\times{\rm GU}_{1}(q)$, then as discussed in Section 3.2, we may assume that the lift $\hat{x}$ of $x$ in ${\rm Sp}_{6}(q)$ is $(\Lambda_{1})^{2},\Lambda_{2}]$ for some $\sigma$-orbits $\Lambda_{1},\Lambda_{2}$. Now we find that $\hat{y}=[\hat{y_{1}},\hat{y_{2}},\hat{y_{3}}]\in\hat{x}^{{\rm Sp}_{2}(q)^{3}}$ if and only if $\hat{y}_{i}$ is ${\rm Sp}_{2}(q)$-conjugate to $\Lambda_{2}]$ for some $i$ and $\hat{y}_{j}$ is ${\rm Sp}_{2}(q)$-conjugate to $\Lambda_{1}$ for $j\in\{1,2,3\}\setminus\{i\}$. Therefore,
@@
Now suppose that $x\in H$ is unipotent. Since $x^{G}\cap H\subseteq B$, we deduce that $x$ must be a $b_{1},c_{2}$ or $b_{3}$ involution by consulting [22, 3.4.14]. We consider each possibility in turn by inspecting [22, Table B.7] and we obtain the bounds in the following Table 7, which allow us to verify that the desired claim holds.
@@
Moreover, with respect to the symplectic basis $\{e_{1},e_{1}^{q},e_{1}^{q^{2}},f_{1},f_{1}^{q},f_{1}^{q^{2}}\}$, $x$ embeds in ${\rm Sp}_{6}(q)$ as $[A,A]$, where $A$ is the matrix defined above. By arguing as in the case $H_{0}={\rm Sp}_{2}(q)\wr S_{3}$, we get
@@
First suppose that $i=6$, so $i_{0}=2$. In this case, $x$ is $G_{I}$-conjugate to $\Lambda]Z$ for some $\sigma$-orbit $\Lambda$, which as discussed in [22, Section 5.3.1] is a union of $3$ distinct $\sigma_{0}$-orbits, where $\sigma_{0}:\alpha\mapsto\alpha^{q^{3}}$. Therefore, $C_{G}(x)$ is of type ${\rm GU}_{1}(q^{3})$ we compute
@@
and we deduce that ${\rm fpr}(x,G/H)<|x^{G}|^{-0.4}$ as desired.
@@
Now assume that $x$ is a field automorphism of order $r$. If $r\neq k$, then [44, 4.9.1] implies that
@@
suffice. Otherwise, if $r=k=2$, then [42, 4.3] gives us
@@
Now assume that $H$ is of type ${\rm GL}_{1}^{\epsilon}(q)\wr S_{5}$. Then $H$ preserves a decomposition $V_{1}\oplus\cdots\oplus V_{5}$ of the natural module $V$. Recall from Lemma 3.18 that $|x^{G_{0}}|>\frac{1}{2}\left(\frac{q}{q+1}\right)^{2}q^{14}$, unless $x$ is semisimple and $C_{G}(x)$ is of type ${\rm GL}_{s}(q)\times{\rm GL}_{n-s}(q)$ for $s\in\{3,4\}$, or $x$ is unipotent with Jordan form $J_{2}^{s},J_{1}^{5-2s}]$ for $s\in\{1,2\}$, and so if $x$ is not in those classes of prime order elements, the trivial bound
$|x^{G}\cap H|\leqslant|H|<\left(\frac{1}{2}\left(\frac{q}{q+1}\right)^{2}q^{14}\right)^{11/20}<|x^{G_{0}}|^{11/20}$ (2)
holds and we are done. We may now assume that $|x^{G_{0}}|\leqslant\frac{1}{2}\left(\frac{q}{q+1}\right)^{2}q^{14}$. Recall that [47, 4.2.9] implies that ${\rm PGL}(V)\cap H\leqslant(q-\epsilon)^{4}.S_{5}=B.S_{5}$.
First assume that $x$ is unipotent. Then $x^{G_{I}}\cap B=\emptyset$, since $r$ does not divide $|B|$, so we may write $x=b\pi$ for some $b\in B$ and $\pi=(r^{h},1^{5-hr})\in S_{5}$. Note that [15, 3.11] implies that $x$ lifts to a unique element $\hat{x}=(\hat{x}_{1},\ldots,\hat{x}_{5})\pi\in\hat{B}\pi$ of order $r=p$, where $\widehat{B}={\rm GL}^{\epsilon}_{1}(q)^{5}$. Moreover, it is shown in the proof of [63, 4.5] that $\hat{x}$ is $\widehat{B}-$conjugate to $(1,\ldots,1,\hat{x}_{hr+1},\ldots,\hat{x}_{5})\pi$, and as demonstrated in the proof of [16, 2.6], $\hat{x}$ has Jordan form $J_{p}^{h+b_{p}},J_{p-1}^{b_{p-1}},\ldots,J_{1}^{b_{1}}]$, where $J_{p}^{b_{p}},\ldots,J_{1}^{b_{1}}]$ is the Jordan form of the restriction of $x$ on $V_{hr+1}\oplus\cdots\oplus V_{5}$.
@@
Now assume that $x$ is semisimple. Recall that under our assumption $C_{G}(x)$ is of type ${\rm GL}_{4}(q)\times{\rm GL}_{1}(q)$ or ${\rm GL}^{\epsilon}_{3}(q)\times{\rm GL}^{\epsilon}_{2}(q)$. First assume that $C_{G}(x)$ is of type ${\rm GL}_{4}(q)\times{\rm GL}_{1}(q)$, so $\hat{x}$ has eigenvalues $(\lambda)^{4},\mu\in\mathbb{F}_{q}$. If $x^{G_{I}}\cap H\subseteq B$, then $|x^{G}\cap H|$ is bounded above by the number of ways of arranging the eigenvalues of $\hat{x}$ on the main diagonal, so we get $|x^{G_{I}}\cap H|\leqslant 5$, and the bound quickly follows, so we may now assume that $x^{G_{I}}\cap(H\setminus B)\neq\emptyset$. We may gain write $x=b\pi$ for some $b\in B$ and $\pi=(r^{h},1^{5-hr})\in S_{5}$. Note that $\pi$ embeds in ${\rm GL}(V)$ as a permutation matrix, and so we quickly deduce that if $\hat{x}=[\lambda I_{4},\mu]$, then $\pi$ must be a transposition, and consequently, $x$ must embed in $G$ as a $t_{1}$ involution. Therefore, we get
$|x^{G_{I}}\cap H|\leqslant|x^{G_{I}}\cap B|+|x^{G_{I}}\cap(H\setminus B)|=5+10(q-\epsilon)$
@@
If $H\in\mathscr{S}$, then as in the proofs of Propositions 4.9, and 4.10, we use the Magma command AutomorphismGroupSimpleGroup() to construct ${\rm Aut}(H)$ and compute $m=\max\{i_{s}(H)\,:\,s\,\text{prime}\}$. If $H$ is of type ${\rm L}_{2}(7)$, then
@@
and this bound is sufficient for all $x\in H$, as $|x^{G_{0}}|>\frac{1}{2}\left(\frac{q}{q+1}\right)q^{4}$ by Lemma 3.18 for all $x\in G_{I}$. On the other hand, if $x\in H$ is of type $A_{6}$, then $m=144$, and since $q\geqslant 23$, we deduce that
@@
If $H$ is of type ${\rm GL}_{1}^{\epsilon}(q^{3})$, then it follows immediately from [22, 5.3.2] that $|x^{G_{0}}|>\frac{1}{6}\left(\frac{q}{q+1}\right)^{2}q^{6}$ for every $x\in H$, and so the trivial bound $|x^{G_{I}}\cap H|\leqslant|H\cap G_{I}|$ is sufficient.
Finally, if $H$ is of type ${\rm GL}^{\epsilon^{\prime}}_{3}(q_{0})$ or ${\rm O}_{3}(q)$, then we obtain the required result by computing exact fixed point ratios using the relevant tables in [22, Appendix B]. We will only consider the case where $\epsilon=+$ and $H$ is of type ${\rm GU}_{3}(q^{1/2})$ here to avoid unnecessary repetition, as all cases are very similar.
First suppose that $x$ is unipotent. Then $x$ either has Jordan form $J_{3}]$ or $J_{2},J_{1}]$ on $V$. If $x$ has Jordan form $J_{3}]$, then we find
@@
If $H$ is of type ${\rm GL}_{3}^{\epsilon^{\prime}}(q_{0})$ or ${\rm O}_{3}(q)$ and $x$ is not an involutory graph-field automorphism then the bounds in the proofs of [15, 5.1] and [15, 8.2] are sufficient, so now assume that $x$ is an involutory graph-field automorphism. In both cases, Lemma 3.8 implies that $x^{G}\cap H\subseteq\widetilde{H}x$, where $\widetilde{H}=H\cap G_{I}$ and using [51, 1.3] we find
$|x^{G}\cap H|\leqslant i_{2}(\widetilde{H})+1<2(1+1q^{-1/2})q^{5/2}$
@@
$|x^{G}\cap H|\subseteq\widetilde{H}x\leqslant i_{2}(\widetilde{H})+1\leqslant 16,$

### v3

@@ 4. Fixed point ratios for classical groups
where $\phi$ is a field automorphism of order 2. We first consider involutions in $x^{G_{I}}\cap B$. Note that, as explained in [22, 3.4.2], $B$ contains two classes of involutions, namely those with centraliser of type ${\rm Sp}_{2}(q^{2})^{2}$ and those with centraliser of type ${\rm GL}_{2}(q^{2})$. If $y$ is of type ${\rm Sp}_{2}(q^{2})^{2}$, then it clearly embeds in $G$ as an involution of type ${\rm Sp}_{4}(q)^{2}$, so $y\in x^{G_{I}}\cap B$. On the other hand, if $y$ is of type ${\rm GL}_{2}(q^{2})$, then the lift $\hat{y}$ of $y$ in ${\rm Sp}_{4}(q^{2})$ has order $4$. Now if $\hat{z}$ denotes the embedding of $\hat{y}$ inside ${\rm Sp}_{8}(q)$, then inspecting [22, Construction 2.4.2] reveals that $\hat{z}$ also has order $4$, so $z=\hat{z}Z({\rm Sp}_{8}(q))$ is of type ${\rm GL}_{4}(q)$. In particular, $y$ embeds in $G_{I}$ as an involution of type ${\rm GL}_{4}(q)$, and so $y\not\in x^{G_{I}}\cap B$. Therefore, we compute
@@
Case 4. $x$ is a field automorphism: In this case, the bounds presented at the start of [16, Section 3.2] and [16, 2.3] are sufficient. ∎
@@
If $x$ is a field, graph, or graph-field automorphism, then the bounds in [15, 8.1] suffice, so it is left to check the claim for semisimple and unipotent elements. In those cases we essentially follow the same strategy as in [15, 8.1]. In particular, we note that $H_{I}=H\cap{\rm PGL}(V)={\rm GSp}_{8}(q)$ and the conjugacy class structure of prime order elements in is briefly discussed in 3.2 and in [22, Chapter 3] in more detail. By using this information and reading off the centraliser structure of prime order elements in ${\rm PGL}_{8}(q)$ and ${\rm PGSp}_{8}(q)$ from [22, Tables B.3 and B.7], we can verify that the claim holds.
For example, suppose that $x$ is a unipotent involution with Jordan decomposition $J_{2}^{2},J_{1}^{4}]$. Then we observe that $x^{G_{I}}\cap H_{I}$ is the union of two $H_{I}$-classes, namely the classes of type $a_{2}$ and $c_{2}$, and so by inspecting [22, Table B.7] we find
$|x^{G_{I}}\cap H|\leqslant\frac{|{\rm Sp}_{8}(q)|}{q^{11}|{\rm Sp}_{2}(q)||{\rm Sp}_{4}(q)|}+\frac{|{\rm Sp}_{8}(q)|}{q^{12}|{\rm Sp}_{4}(q)|}=(q^{8}-1)(q^{6}+q^{4}+q^{2})$
and
$|x^{G_{0}}|\geqslant\frac{|{\rm GL}_{8}(q)|}{q^{9}|{\rm GL}_{2}(q)||{\rm GL}_{4}(q)|}\geqslant q^{23}(q^{5}-1)(q^{7}-1)$
so the claim holds. We deal with all other cases in a similar fashion. ∎
@@
    $\epsilon=+$, $x$ is semisimple with $C_{G}(x)$ of type ${\rm GL}_{3}(q)^{2}$, and $\mbox{{\rm fpr}}(x,G/H)<|x^{G_{0}}|^{-0.32}$
@@
    $\epsilon=$-, $x$ is semisimple with $C_{G}(x)$ of type ${\rm GL}_{3}(q^{2})$, and $\mbox{{\rm fpr}}(x,G/H)<|x^{G_{0}}|^{-0.32}$.
@@
Here we may assume that $H$ is of type ${\rm Sp}_{6}(q)$ since the claim follows directly by Theorem 3.9 in all other cases.
Case 1. $x$ is semisimple and $r>2$: Let $i$ be minimal such that $r|q^{i}-1$. Again, we may assume that $x\in H$ and in particular that $r$ divides $|H|$, since otherwise we have ${\rm fpr}(x,G/H)=0$. Thus, we have $i\in\{1,2,3,4,6\}$. The embedding here is straightforward so we will only give details for the cases where $i\in\{1,4\}$ and leave the remaining cases for the reader.
@@
Case 2. $x$ is unipotent: By appealing to Section 3.2, we find that the Jordan decompositions corresponding to an element $x\in H$ are
@@
Case 3. $x$ is a graph, field or graph-field automorphism or a semisimple involution: Here the bounds in the proof of [15, 8.1] suffice. ∎
Remark 4.7.
The cases in (i) and (ii) of Proposition 4.6 are genuine exceptions. In particular, the bounds we have derived are precise and one can check that the functions
$f(x)=\frac{\log(x^{6}(x+1)(x^{2}+1)(x^{3}+1))}{\log(x^{9}(x^{4}+x^{3}+x^{2}+x+1)(x^{2}+1)(x^{3}+1))}$
and
$g(x)=\frac{\log(x^{6}(x+1)(x^{2}+1)(x^{3}+1))}{\log(x^{9}(x+1)(x^{3}+1)(x^{5}+1))}$
both are decreasing on $(0,\infty)$ with a horizontal asymptote at $2/3$, so we cannot bound ${\rm fpr}(x,G/H)$ by $|x^{G_{0}}|^{-1/3}$.
@@
The claim can be easily verified using Magma for $q\leqslant 4$, and so we may assume that $q\geqslant 8$ for the remainder of the proof. By inspecting [12, Tables 8.28 and 8.29], we find that if $H$ is not of type $G_{2}(q)$, then $H_{0}=H\cap G_{0}\in\{{\rm Sp}_{2}(q)\wr S_{3},{\rm Sp}_{2}(q^{3}){:}3,{\rm Sp}_{6}(q_{0})\}$, where $q=q_{0}^{k}$ for some prime $k$.
@@
Now suppose that $x\in{\rm PGL}(V)$. There are two cases to consider depending whether or not $x^{G}\cap(H\setminus B)\neq\emptyset$. First suppose that $x^{G}\cap(H\setminus B)\neq\emptyset$. In this case, we note that either $r=2$ or $r=3$. If $r=2$, then by inspecting the proof of [16, 2.8], we deduce that $\log|x^{G_{I}}\cap H|/\log|x^{G_{0}}|<0.543$, so we may now assume that $r=3$. Here, we may assume $x=b\pi=(x_{1},x_{2},x_{3})\pi$, where $b\in B$ and $\pi\in S_{3}$ is a $3$-cycle. It then follows from the proof of [64, 4.5] that $x$ is $B$-conjugate to $\pi$. Now note that if $y=(X,Y,Z)\in C_{B}(x)$, then since $\pi$ permutes $X,Y$, and $Z$, we must have $X=Y=Z$, and so
@@
We now note that with respect to a symplectic basis $\{e_{1},e_{2},e_{3},f_{1},f_{2},f_{3}\}$, $\pi$ embeds in ${\rm Sp}_{6}(q)$ as the block-diagonal matrix $[A,A]$, where
$A=\begin{pmatrix}0&0&1\\ 1&0&0\\ 0&1&0\end{pmatrix}$. (2)
Since $r=3$ and $q$ is even, it follows that $x$ is semisimple. Moreover, since $q$ is even, $3$ must either divide q1q-1 or $q+1$, and so if $i$ is minimal such that $3$ divides $q^{i}-1$, we must have $i=1$ or $2$.
Now note that $x$ has eigenvalues $1^{2},(\omega)^{2},(\omega^{2})^{2}$, where $\omega\in\mathbb{F}_{q^{i}}$ is a primitive cube root of unity, and so $C_{G}(x)$ is of type ${\rm GL}_{2}^{\epsilon}(q)\times{\rm Sp}_{2}(q)$, where $\epsilon=(-1)^{i+1}$. Therefore,
@@
We can now compute bounds for $|x^{G_{I}}\cap H|$ and $|x^{G_{0}}|$ using [22, Table B.7]. For example, if $C_{G}(x)$ is of type ${\rm GU}_{2}(q)\times{\rm GU}_{1}(q)$, then as discussed in Section 3.2, we may assume that the lift $\hat{x}$ of $x$ in ${\rm Sp}_{6}(q)$ is $(\Lambda_{1})^{2},\Lambda_{2}]$ for some $\sigma$-orbits $\Lambda_{1},\Lambda_{2}$. Now we find that $\hat{y}=[Y_{1},Y_{2},Y_{3}]\in\hat{x}^{{\rm Sp}_{2}(q)^{3}}$ if and only if $Y_{i}$ is ${\rm Sp}_{2}(q)$-conjugate to $\Lambda_{2}]$ for some $i$ and $Y_{j}$ is ${\rm Sp}_{2}(q)$-conjugate to $\Lambda_{1}$ for $j\neq i$. Therefore,
@@
Now suppose that $x\in H$ is unipotent. Since $x^{G}\cap H\subseteq B$, we deduce that $x$ must be a $b_{1},c_{2}$ or $b_{3}$ involution by consulting [22, 3.4.14]. We consider each possibility in turn by inspecting [22, Table B.7] and we obtain the bounds in Table 8, which allow us to verify that the desired claim holds.
@@
Moreover, with respect to the symplectic basis $\{e_{1},e_{1}^{q},e_{1}^{q^{2}},f_{1},f_{1}^{q},f_{1}^{q^{2}}\}$, $x$ embeds in ${\rm Sp}_{6}(q)$ as $[A,A]$, where $A$ is the matrix defined in (2). By arguing as in the case $H_{0}={\rm Sp}_{2}(q)\wr S_{3}$, we get
@@
First suppose that $i=6$, so $i_{0}=2$. In this case, $x$ is $G_{I}$-conjugate to $\Lambda]Z$ for some $\sigma$-orbit $\Lambda$, which as discussed in [22, Section 5.3.1] is a union of $3$ distinct $\sigma_{0}$-orbits, where $\sigma_{0}:\alpha\mapsto\alpha^{q^{3}}$. Therefore, $C_{G}(x)$ is of type ${\rm GU}_{1}(q^{3})$, and we compute
@@
and we deduce that ${\rm fpr}(x,G/H)<|x^{G_{0}}|^{-0.4}$ as desired.
@@
Now assume that $x$ is a field automorphism of order $r$. If $r\neq k$, then [45, 4.9.1] implies that
@@
suffice. Otherwise, if $r=k=2$, then [43, 4.3] gives us
@@
Now assume that $H$ is of type ${\rm GL}_{1}^{\epsilon}(q)\wr S_{5}$. Then $H$ preserves a decomposition $V_{1}\oplus\cdots\oplus V_{5}$ of the natural module $V$. Recall from Lemma 3.18 that $|x^{G_{0}}|>\frac{1}{2}\left(\frac{q}{q+1}\right)^{2}q^{14}$, unless $x$ is semisimple and $C_{G}(x)$ is of type ${\rm GL}^{\epsilon}_{s}(q)\times{\rm GL}^{\epsilon}_{n-s}(q)$ for $s\in\{3,4\}$, or $x$ is unipotent with Jordan form $J_{2}^{s},J_{1}^{5-2s}]$ for $s\in\{1,2\}$. Therefore, if $x$ is not in those classes of prime order elements, we may use the trivial bound
$|x^{G}\cap H|\leqslant|H|<\left(\frac{1}{2}\left(\frac{q}{q+1}\right)^{2}q^{14}\right)^{11/20}<|x^{G_{0}}|^{11/20}.$ (4)
We may now assume that $|x^{G_{0}}|\leqslant\frac{1}{2}\left(\frac{q}{q+1}\right)^{2}q^{14}$. Recall that [48, 4.2.9] implies that ${\rm PGL}(V)\cap H\leqslant(q-\epsilon)^{4}.S_{5}=B.S_{5}$. First assume that $x$ is unipotent. Then $x^{G_{I}}\cap B=\emptyset$, since $r$ does not divide $|B|$, so we may write $x=b\pi$ for some $b\in B$ and $\pi=(r^{h},1^{5-hr})\in S_{5}$. Note that [15, 3.11] implies that $x$ lifts to a unique element $\hat{x}=(\hat{x}_{1},\ldots,\hat{x}_{5})\pi\in\hat{B}\pi$ of order $r=p$, where $\widehat{B}={\rm GL}^{\epsilon}_{1}(q)^{5}$. Moreover, it is shown in the proof of [64, 4.5] that $\hat{x}$ is $\widehat{B}-$conjugate to $(1,\ldots,1,\hat{x}_{hr+1},\ldots,\hat{x}_{5})\pi$, and as demonstrated in the proof of [16, 2.6], $\hat{x}$ has Jordan form $J_{p}^{h+b_{p}},J_{p-1}^{b_{p-1}},\ldots,J_{1}^{b_{1}}]$, where $J_{p}^{b_{p}},\ldots,J_{1}^{b_{1}}]$ is the Jordan form of the restriction of $x$ on $V_{hr+1}\oplus\cdots\oplus V_{5}$.
@@
Now assume that $x$ is semisimple. Recall that under our assumption $C_{G}(x)$ is of type ${\rm GL}^{\epsilon}_{4}(q)\times{\rm GL}^{\epsilon}_{1}(q)$ or ${\rm GL}^{\epsilon}_{3}(q)\times{\rm GL}^{\epsilon}_{2}(q)$. First assume that $C_{G}(x)$ is of type ${\rm GL}^{\epsilon}_{4}(q)\times{\rm GL}^{\epsilon}_{1}(q)$, so $\hat{x}$ has eigenvalues $(\lambda)^{4},\mu\in\mathbb{F}_{q}$. If $x^{G_{I}}\cap H\subseteq B$, then $|x^{G}\cap H|$ is bounded above by the number of ways of arranging the eigenvalues of $\hat{x}$ on the main diagonal, so we get $|x^{G_{I}}\cap H|\leqslant 5$, and the bound quickly follows. We may now assume that $x^{G_{I}}\cap(H\setminus B)\neq\emptyset$. We may gain write $x=b\pi$ for some $b\in B$ and $\pi=(r^{h},1^{5-hr})\in S_{5}$. Note that $\pi$ embeds in ${\rm GL}(V)$ as a permutation matrix, and so we quickly deduce that if $\hat{x}=[\lambda I_{4},\mu]$, then $\pi$ must be a transposition, and consequently, $x$ must embed in $G$ as a $t_{1}$-involution. Therefore, we get
$|x^{G_{I}}\cap H|=|x^{G_{I}}\cap B|+|x^{G_{I}}\cap(H\setminus B)|\leqslant 5+10(q-\epsilon)$
@@
If $H\in\mathscr{S}$, then as in the proofs of Propositions 4.10, and 4.11, we use the Magma command AutomorphismGroupSimpleGroup() to construct ${\rm Aut}(H)$ and compute $m=\max\{i_{s}(H)\,:\,s\,\text{prime}\}$. If $H$ is of type ${\rm L}_{2}(7)$, then
@@
and this bound is sufficient for all $x\in H$, as $|x^{G_{0}}|>\frac{1}{2}\left(\frac{q}{q+1}\right)q^{4}$ by Lemma 3.18. On the other hand, if $H$ is of type $A_{6}$, then $m=144$, and since $q\geqslant 23$, we deduce that
@@
If $H$ is of type ${\rm GL}_{1}^{\epsilon}(q^{3})$, then it follows immediately from [22, 5.3.2] that $|x^{G_{0}}|>\frac{1}{6}\left(\frac{q}{q+1}\right)^{2}q^{6}$, and so the trivial bound $|x^{G_{I}}\cap H|\leqslant|H\cap G_{I}|$ is sufficient.
Finally, if $H$ is of type ${\rm GL}^{\epsilon^{\prime}}_{3}(q_{0})$ or ${\rm O}_{3}(q)$, then we obtain the required result by computing exact fixed point ratios using the relevant tables in [22, Appendix B]. All cases are very similar, so we will only consider the case where $\epsilon=+$ and $H$ is of type ${\rm GU}_{3}(q^{1/2})$ here to avoid unnecessary repetition.
First suppose that $x$ is unipotent. Then $x$ has Jordan form $J_{3}]$ or $J_{2},J_{1}]$ on $V$. If $x$ has Jordan form $J_{3}]$, then we find
@@
If $H$ is of type ${\rm GL}_{3}^{\epsilon^{\prime}}(q_{0})$ or ${\rm O}_{3}(q)$ and $x$ is not an involutory graph-field automorphism then the bounds in the proofs of [15, 5.1] and [15, 8.2] are sufficient, so now assume that $x$ is an involutory graph-field automorphism. In both cases, Lemma 3.8 implies that $x^{G}\cap H\subseteq H_{I}x$, where $H_{I}=H\cap G_{I}$ and using [52, 1.3] we find
$|x^{G}\cap H|\leqslant i_{2}(H_{I})+1<2(1+1q^{-1/2})q^{5/2}$
@@
$|x^{G}\cap H|\subseteq H_{I}x\leqslant i_{2}(H_{I})+1\leqslant 16,$